A jar with pennies, nickels, and dimes contains a total of 55 coins worth $3.23. How many of each are there if the number of pennies is 10 less than the sum of nickels and dimes together?
Answer by ptaylor(2198)

YOU NEED TO RECHECK YOUR PROBLEM. UNLESS I MADE A MISTAKE, WE ARE COMING OUT WITH FRACTIONS OF COINS AND WE CAN'T HAVE FRACTIONS OF COINS.
Let x=number of pennies
y=number of nickels
z=number of dimes
Then:
x+y+z=55-----------------eq1
Lets deal in pennies
x+5y+10z=323------------------eq2
and
x+10=y+z or
x-y-z=-10 or
x-(y+z)=-10----------------------eq3
but from eq1, y+z=55-x. Substitute this into eq3
x-(55-x)=-10 or
x-55+x=-10 add 55 to each side
2x=45-----------------------------------------NO GOOD!!!!
LET'S TRY ANOTHER APPROACH:
subtract eq1 from eq2 and we get:
4y+9z=268-----------------------------------eq1a
subtract eq3 from eq2 and we get:
6y+11z=333------------------------------------eq2a
Multiply eq1a by 3:
12y+27z=804----------------------eq1b
Multiply eq2a by 2:
12y+22z=666--------------------------eq2b
subtract eq2b from eq1b:
5z=138-------------------------------------------NO GOOD!!!!!
Hope this helps---ptaylor